Suppose 2.17 mol of HCl is dissolved in enough water to give 504.5 mL of solution.
What is the H3O+ concentration?
What is the pH?
1
Expert's answer
2020-05-10T09:20:40-0400

HCl is a strong electrolyte: it means that the dissociation is nearly complete in water solution (pKa=-9.3, or Ka=109.3). The reaction of dissociation of HCl in water is:

HCl + H2O "\rightarrow" H3O+ + Cl-.

As you can see, one mole of HCl gives one mole of the H3O+ ions. Therefore, the concentration of H3O+ is equal to the concentration of HCl:

"c(H_3O^+) = c(HCl) = \frac{2.17}{504.5·10^{-3}} = 4.30" mol/L, or M.

The pH is defined as -log(c(H+)):

"pH = -log([H_3O^+]) = -log(4.30) = -0.63" .

Don't worry about the negative pH. The misconception that pH lies between 0 and 14 is disclaimed in literature[1].


[1] : Kieran F. Lim, Journal of Chemical Education 2006 83 (10), 1465, DOI: 10.1021/ed083p1465

Answer: the concentration of H3O+ is 4.30 M. The pH is -0.63.